OpenStudy (anonymous):

How many times more acidic is acid rain with a pH of 2.4 than an ordinary rain with pH of 6?

sam (.sam.):

Moving from pH of 6 to 5 is about 10 times stronger.

OpenStudy (anonymous):

May I ask what formula did you use, if you used any at all?

sam (.sam.):

-log[H+]=pH

OpenStudy (anonymous):

thats odd my textbook says the formula is [H+]=10^pH

OpenStudy (anonymous):

*-pH

sam (.sam.):

its the same, you take the log of both sides

sam (.sam.):

-log[H+]=log10^pH -log[H+]=pHlog10 -log[H+]=pH

OpenStudy (anonymous):

so if pH is 2.4....damn hold on a sec...

OpenStudy (anonymous):

oh i see, so its (10^-2.4/10^6)=10^4.6?

sam (.sam.):

yes

OpenStudy (anonymous):

and I can just state it was ten times stronger, not 10,000?

sam (.sam.):

no don't say 10,000 times stronger that will be inaccurate

OpenStudy (anonymous):

alright I'll stick with 10 times stronger thank you so much I have a final and I'm trying to study hard right now. Have a good night.

sam (.sam.):

NOOO, don't use 10 times! that's just a definition, The answer is -log(3.98x10^-3)=2.4 -log(1x10^-6)=6 Divide both concentrations, 3.98x10^-3 / 1x10^-6 You will get 3980 time stronger